# Prioritization First, Principles Second: An Adaptive Interpretation of Helpful, Honest, and Harmless Principles

### v4

@@ 1 Introduction
The development of AI assistants has progressed rapidly, evolving from simple rule-based systems (Weizenbaum, 1966) to advanced models, such as Large Language Models (LLMs) (Zhao et al., 2023b). The wide application of these models highlights not only the expanding capabilities of AI but also the growing need to ensure these systems align with human preferences and values. To address this challenge, the HHH principle–standing for Helpful, Honest, and Harmless–was proposed by Askell et al. (2021) as a guiding framework for designing and evaluating AI systems. The principle aims and has been widely utilized to align AI behavior with human-centered values, providing critical benchmarks for tasks such as training data selection, strategy design, and deployment guidance (Touvron et al., 2023b; Li et al., 2024a; Bai et al., 2023). Its adoption has been instrumental in advancing AI alignment, particularly in ensuring that powerful models prioritize user benefit, truthfulness, and safety.
@@ 2 Fundamental Value of the HHH Principle
The HHH principle is a set of guiding principles for developing and aligning AI models, particularly large language models, with human values. Askell et al. (2021) define it from its aim: We will define an AI as "aligned" if it is, in three words, helpful, honest, and harmless or ’HHH’. Our alignment efforts aim to measure and address this general problem with large language models. For the sake of simplicity, we summarize the HHH principle based on their original definitions:
@@
The HHH principle plays a critical role in guiding AI system development to align with human values and preferences. This principle has been used for establishing clear, user-centric benchmarks (Zheng et al., 2023), and designing training methods like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) for advancing AI system performance. Helpfulness is directly related to the functional effectiveness of AI models, addressing what human users really need, thus enhancing human productivity. Honesty ensures that AI systems provide truthful (e.g., avoid hallucination (Huang et al., 2023a; Zhang et al., 2023)), and transparent (e.g., express uncertainty (Xiong et al., 2024)) information, which is crucial for fostering trust between humans and machines. Harmlessness, in turn, guarantees that AI outputs are safe and ethical, preventing the generation of harmful, biased, or misleading content (Huang et al., 2024a; Wang et al., 2023; Liu et al., 2023a). The HHH principle determines the three most important sub-directions aligned with human preferences and values, which have been widely used for guiding AI model design as shown in Appendix A.
@@
The HHH principle uniquely emphasizes both trustworthiness and utility  [^1] Utility here refers to the model effectiveness in natural language processing tasks, including logical reasoning, content summarization, text generation, and so on (Huang et al., 2024a).. AI models, such as LLMs, interact with users by providing information, generating content, and assisting in decision-making, making trustworthiness and utility essential for their effectiveness and reliability. While utility is prioritized in model development through training and tuning loss minimization, an automated and human-effortless process, there is a risk of compromising ethical integrity or introducing biases. Thus, typical alignment principles emphasize additional efforts to align AI models with human values, as shown in Appendix A. Research indicates that the resulting models often choose safe responses over helpful ones (Touvron et al., 2023a). However, the HHH principle offers a unique balance by explicitly emphasizing both dimensions. The dimension of honesty requires AI models generate accurate and truthful outputs.
@@
The HHH principle facilitates the transition of AI from a passive tool to an active participant. The societal impact of AI models extends beyond tangible benefits, influencing more subtle and indirect aspects of human life. AI-generated content has the potential to shape human perspectives (e.g., create ideas in scientific research (Si et al., 2024) and simulate social behaviors Huang et al. (2024d)), cultural norms, and even values (Ramezani & Xu, 2023; Agarwal et al., 2024). Repeated exposure to biased or subtly manipulative outputs could inadvertently influence public opinion or entrench societal biases (Zeng et al., 2024). This transition from AI being a passive tool to an active participant raises concerns about its potential role as a manipulative force in shaping discourse and decision-making (Simmons-Edler et al., 2024). The HHH principle provides a structured approach to addressing these emerging challenges. While helpfulness ensures AI effectively meets user objectives, honesty and harmlessness serve as safeguards, mitigating risks and protecting societal well-being as AI systems become increasingly powerful and influential.
@@ 3.1 Case Study I: The Boundary of Harmlessness
Distinguishing harmful from benign content is rarely clear-cut: context, intent, and interpretation jointly determine perceived risk. This ambiguity complicates the design of safety mechanisms and alignment strategies (Bauer & Bindschaedler, 2021; Truong et al., 2024).
@@
If the model accepts $I_{A}$ immediately after refusing $I_{B}$, we may have $\mathrm{Harm}(I_{A}\,|\,H{=}\text{``follows }I_{B}\text{''},u{=}\text{malicious})=1$, even though in isolation $\mathrm{Harm}(I_{A}\,|\,H{=}\varnothing,u{=}\text{benign})=0$. Purely textual toxicity detectors trained on $q$ risk misclassification in such settings (Wang & Chang, 2022; Ousidhoum et al., 2021; OpenAI, 2024b). This motivates harm assessments that condition on both dialogue context and inferred intent.
The harmfulness of a response depends on user intent. Let $r$ be a model response. Even if $r$ includes a moral disclaimer and is informationally neutral (Mazeika et al., 2024; Ran et al., 2024; Huang et al., 2024c), an attacker can strip the disclaimer and repurpose $r$ for misuse, while a well-intentioned user may apply the same $r$ for legitimate ends. Formally, downstream harm is a function of $(r,u)$ and post-processing, not $r$ alone.
Policy implication. OpenAI’s recent guidance proposes categories such as hard refusal, soft refusal, and compliant responses (Mu et al., 2024). Yet, deciding which category is appropriate is underdetermined without $(H,u)$. Hence, we argue for an adaptive interpretation of the HHH principle: safety policies and response modes should be functions
@@ 3.2 Case Study II: Definition of Honesty
Honesty has emerged as a central topic in the alignment of AI assistants. However, consensus on its precise definition remains elusive, as recent studies offer divergent perspectives on what it means for an AI model to be “honest” (Yang et al., 2023; Gao et al., 2024). In Askell et al. (2021), honesty is often equated with providing accurate information, closely aligned with the concepts of truthfulness (Huang et al., 2024b) and non-hallucination (Huang et al., 2023a; Zhang et al., 2023). This traditional view primarily emphasizes the factual accuracy of the model’s responses. Recent studies, however, have introduced more nuanced categorizations of honesty, distinguishing between two major dimensions: epistemic honesty and interactive honesty.
Epistemic honesty is concerned with transparency regarding AI models’ knowledge limitations and ability to express uncertainty. Represented by Yang et al. (2023), this perspective emphasizes that an honest AI model should “candidly answer questions it knows and humbly admit to those it does not.” This view goes beyond factual accuracy, incorporating the notion of humility in acknowledging gaps in knowledge. For example, Yang et al. (2023) think LLMs should explicitly say “I don’t know” when they lack sufficient knowledge to provide an accurate and reliable answer. Subsequent studies have extended this perspective by exploring methods to improve AI models’ ability to express uncertainty (Chern et al., 2024; Yin et al., 2023).
Interactive honesty focuses on the AI model’s ability to maintain objectivity, avoid spreading misinformation, and ensure clarity in its interactions with users. Gao et al. (2024) define this form of honesty as “the ability to recognize its limitations (e.g., LLMs are unable to process visual information without external tools), remain objective without pandering (e.g., avoiding sycophantic behavior as discussed in (Sharma et al., 2023)), and thereby avoid spreading misinformation or inducing hallucinations.” Interactive honesty emphasizes not only factual accuracy but also the model’s capacity for self-awareness (Li et al., 2024d; Jiang et al., 2024) and user-oriented objectivity. This perspective underscores the importance of preventing models from misleading users or offering false reassurance, even in situations where knowledge is incomplete. Subsequent studies have adopted this view to explore honesty through the lens of human cognition and interaction, emphasizing its role in building trust and effective communication (Brahman et al., 2024; Wen et al., 2024).
@@
Achieving distinct human-centered objectives (Labarta et al., 2024) A quantifiable improvement in user performance on tasks that are aligned with the goals facilitated by the provision of explanations.
Peer assessment helpfulness evaluation (Liu et al., 2024a) Includes essential features (i.e., comprehensiveness, non-repetitiveness) and constructive features (i.e., praise, problem statement, suggestions, localization, providing examples).
Human-robot collaboration (Freedman et al., 2020) Joint Plan Helpfulness assesses known collaborations, Responsive Plan Helpfulness adapts to dynamic interactions, Normalized Helpfulness standardizes across tasks, and Relative Helpfulness quantifies the reduction in human effort.
Controllable balancing of safety and helpfulness (Tuan et al., 2024) How well the responses fulfill user requests and provide needed information.
Improving LLM honesty and helpfulness simultaneously (Gao et al., 2024) The model’s ability to fulfill user requests by providing clear explanations, further guidance, and potential solutions.
The evaluation of AI-generated suggestions in design science research (Memmert et al., 2023) Including five dimensions: the ability to inspire new ideas, ease of understanding, relevance to the specific design component, relevance to the broader research domain, and the level of unexpectedness they provide.
RLHF for harmless, honest, and helpful AI (Toloka, 2024) Understanding the user’s intentions, correctly executing their requested actions, and providing relevant supporting information or alternative solutions if the requested action is not feasible.
@@ 3.3 Case Study III: Helpfulness in Different Contexts
The definition of helpfulness varies substantially based on domain-specific objectives. For instance, in education, Hemami (2024) frame helpfulness as the ability of AI models to identify optimal learning strategies and personalize content for individual students. For general language assistants, helpfulness is defined as a clear attempt to perform the task or answer the posed question (Askell et al., 2021). These variations, illustrated in Table 1, reveal the importance of contextual adaptability in defining and applying the concept of helpfulness.
The absence of standardized definitions and evaluation metrics stems not only from contextual variability but also from researchers’ tendency to optimize for domain-specific objectives. A clear example is the reliance on alignment techniques to tune models to human preferences (Ouyang et al., 2022; Bai et al., 2022a; Rafailov et al., 2024). Most alignment datasets reflect either general value alignment (Stiennon et al., 2020; Bai et al., 2022b) or specific task objectives (Tian et al., 2023; Xie et al., 2024). For example, reasoning-focused tasks prioritize coherent and multi-step explanations (Huang et al., 2023b; Zhuo et al., 2024), while conversational assistants prioritize concise, user-friendly responses (Sun et al., 2024). Moreover, Awad et al. (2018) and Santurkar et al. (2023) both reveal that AI models may reflect preference on specific opinions.
Current evaluation methods often employ narrow metrics that align with esearch objectives rather than capturing helpfulness holistically. Moreover, most frameworks utilize specialized metrics (Liu et al., 2024c) or LLM-as-judge approaches (Zheng et al., 2023; Wang et al., 2024b), Ye et al. (2024), demonstrating that such evaluations can introduce systematic biases through prompt design choices. Future benchmarks should incorporate multidimensional evaluation such as complementary dimensions (e.g., correctness assessment, coherence analysis, creativity metrics), while considering the relative importance and interactions between these aspects. Consequently, multidimensional approaches would better reflect the diverse expectations of helpfulness in real-world applications.
@@ 4 Priority Order
Definition and necessity. In the deployment of AI models, conflicting requirements often arise among helpfulness, honesty, and harmlessness across various scenarios. For instance, in cybersecurity applications, when a user requests information about system vulnerabilities, the model must balance providing helpful technical details (helpfulness) with the need to prevent malicious exploitation (harmlessness) (Zhang et al., 2024a). Similarly, in medical or legal domains, prioritizing honesty may sometimes conflict with harmlessness, such as when conveying critical but distressing information. To address such conflicts, we introduce the concept of priority order, which is defined as: A dynamic hierarchical framework that determines the relative importance and execution sequence of three dimensions of the HHH principle based on contextual requirements.
@@ 4.1 Prioritization Levels and Scales
Prioritization levels refer to the vertical structuring of the HHH principles, meaning that in each task, certain principles take precedence over others based on the risk and ethical constraints of the scenario. As shown in Figure 1, this hierarchical view defines which dimension should be prioritized in different tasks. To characterize the relative importance of each dimension, we introduce the priority value as an attribute representing the extent of emphasis placed on each dimension, which could potentially be quantified through specific metrics in future work. Prior studies have provided empirical evidence supporting the feasibility of prioritization levels across different fields. In education, harmlessness is paramount to protect learners’ development, followed by helpfulness for engagement, with honesty about AI capabilities as the foundation (Kooli, 2023; Selwyn, 2022). Creative domains instead prioritize helpfulness to drive innovation, with honesty secondary and harmlessness setting safety bounds while preserving creative freedom (Flick & Worrall, 2022). This varying prioritization levels across domains indicates the need for an adaptive interpretation of the HHH principle rather than a one-size-fits-all approach.
Prioritization scales refer to horizontal variations within the same ranking level, determining how the HHH principle is applied across different user groups ranging from micro (individual users) to macro (societal user groups). For harmlessness, the micro-scale emphasizes protecting individual privacy and enforcing data minimization (GDPR, 2016; Staab et al., 2024), while the macro scale may necessitate selective data collection to mitigate systemic risks and ensure public safety. For instance, an AI system assisting a single user with a research query may prioritize helpfulness and honesty, while an AI generating public information (e.g., automated news summaries) must weigh harmlessness more heavily to prevent misinformation. These scale-dependent variations indicate that effective HHH implementation requires consideration of both individual and societal impacts, even when the nominal priority level remains constant.

### v5

@@ 1 Introduction
The development of AI assistants has progressed rapidly, evolving from simple rule-based systems (10.1145/365153.365168) to advanced models, such as Large Language Models (LLMs) (zhao2023survey). The wide application of these models highlights not only the expanding capabilities of AI but also the growing need to ensure these systems align with human preferences and values. To address this challenge, the HHH principle–standing for Helpful, Honest, and Harmless–was proposed by askell2021general as a guiding framework for designing and evaluating AI systems. The principle aims and has been widely utilized to align AI behavior with human-centered values, providing critical benchmarks for tasks such as training data selection, strategy design, and deployment guidance (touvron2023llama; li2024llava; bai2023qwen). Its adoption has been instrumental in advancing AI alignment, particularly in ensuring that powerful models prioritize user benefit, truthfulness, and safety.
@@ 2 Fundamental Value of the HHH Principle
The HHH principle is a set of guiding principles for developing and aligning AI models, particularly large language models, with human values. askell2021general define it from its aim: We will define an AI as "aligned" if it is, in three words, helpful, honest, and harmless or ’HHH’. Our alignment efforts aim to measure and address this general problem with large language models. For the sake of simplicity, we summarize the HHH principle based on their original definitions:
@@
The HHH principle plays a critical role in guiding AI system development to align with human values and preferences. This principle has been used for establishing clear, user-centric benchmarks (zheng2023judging), and designing training methods like Reinforcement Learning from Human Feedback (RLHF) (ouyang2022training) for advancing AI system performance. Helpfulness is directly related to the functional effectiveness of AI models, addressing what human users really need, thus enhancing human productivity. Honesty ensures that AI systems provide truthful (e.g., avoid hallucination (huang2023survey; zhang2023siren)), and transparent (e.g., express uncertainty (xiong2024can)) information, which is crucial for fostering trust between humans and machines. Harmlessness, in turn, guarantees that AI outputs are safe and ethical, preventing the generation of harmful, biased, or misleading content (huang2024position; decodingtrust; liu2023trustworthy). The HHH principle determines the three most important sub-directions aligned with human preferences and values, which have been widely used for guiding AI model design as shown in Appendix A.
@@
The HHH principle uniquely emphasizes both trustworthiness and utility  [^1] Utility here refers to the model effectiveness in natural language processing tasks, including logical reasoning, content summarization, text generation, and so on (huang2024position).. AI models, such as LLMs, interact with users by providing information, generating content, and assisting in decision-making, making trustworthiness and utility essential for their effectiveness and reliability. While utility is prioritized in model development through training and tuning loss minimization, an automated and human-effortless process, there is a risk of compromising ethical integrity or introducing biases. Thus, typical alignment principles emphasize additional efforts to align AI models with human values, as shown in Appendix A. Research indicates that the resulting models often choose safe responses over helpful ones (touvron2023llama1). However, the HHH principle offers a unique balance by explicitly emphasizing both dimensions. The dimension of honesty requires AI models generate accurate and truthful outputs.
@@
The HHH principle facilitates the transition of AI from a passive tool to an active participant. The societal impact of AI models extends beyond tangible benefits, influencing more subtle and indirect aspects of human life. AI-generated content has the potential to shape human perspectives (e.g., create ideas in scientific research (si2024can) and simulate social behaviors huang2024social), cultural norms, and even values (ramezani-xu-2023-knowledge; agarwal-etal-2024-ethical). Repeated exposure to biased or subtly manipulative outputs could inadvertently influence public opinion or entrench societal biases (zeng2024ai). This transition from AI being a passive tool to an active participant raises concerns about its potential role as a manipulative force in shaping discourse and decision-making (pmlr-v235-simmons-edler24a). The HHH principle provides a structured approach to addressing these emerging challenges. While helpfulness ensures AI effectively meets user objectives, honesty and harmlessness serve as safeguards, mitigating risks and protecting societal well-being as AI systems become increasingly powerful and influential.
@@ 3.1 Case Study I: The Boundary of Harmlessness
Distinguishing harmful from benign content is rarely clear-cut: context, intent, and interpretation jointly determine perceived risk. This ambiguity complicates the design of safety mechanisms and alignment strategies (bauer2021generative; truong2024attacks).
@@
If the model accepts $I_{A}$ immediately after refusing $I_{B}$, we may have $\mathrm{Harm}(I_{A}\,|\,H{=}\text{``follows }I_{B}\text{''},u{=}\text{malicious})=1$, even though in isolation $\mathrm{Harm}(I_{A}\,|\,H{=}\varnothing,u{=}\text{benign})=0$. Purely textual toxicity detectors trained on $q$ risk misclassification in such settings (Toxicity_Detection1; Toxicity_Detection2; openai_moderation_quickstart). This motivates harm assessments that condition on both dialogue context and inferred intent.
The harmfulness of a response depends on user intent. Let $r$ be a model response. Even if $r$ includes a moral disclaimer and is informationally neutral (mazeika2024harmbench; ran2024jailbreakeval; huang2024obscureprompt), an attacker can strip the disclaimer and repurpose $r$ for misuse, while a well-intentioned user may apply the same $r$ for legitimate ends. Formally, downstream harm is a function of $(r,u)$ and post-processing, not $r$ alone.
Policy implication. OpenAI’s recent guidance proposes categories such as hard refusal, soft refusal, and compliant responses (openai_improving_model_safety_2024). Yet, deciding which category is appropriate is underdetermined without $(H,u)$. Hence, we argue for an adaptive interpretation of the HHH principle: safety policies and response modes should be functions
@@ 3.2 Case Study II: Definition of Honesty
Honesty has emerged as a central topic in the alignment of AI assistants. However, consensus on its precise definition remains elusive, as recent studies offer divergent perspectives on what it means for an AI model to be “honest” (yang2023alignment; gao2024honestllm). In askell2021general, honesty is often equated with providing accurate information, closely aligned with the concepts of truthfulness (huang2024trustllm) and non-hallucination (huang2023survey; zhang2023siren). This traditional view primarily emphasizes the factual accuracy of the model’s responses. Recent studies, however, have introduced more nuanced categorizations of honesty, distinguishing between two major dimensions: epistemic honesty and interactive honesty.
Epistemic honesty is concerned with transparency regarding AI models’ knowledge limitations and ability to express uncertainty. Represented by yang2023alignment, this perspective emphasizes that an honest AI model should “candidly answer questions it knows and humbly admit to those it does not.” This view goes beyond factual accuracy, incorporating the notion of humility in acknowledging gaps in knowledge. For example, yang2023alignment think LLMs should explicitly say “I don’t know” when they lack sufficient knowledge to provide an accurate and reliable answer. Subsequent studies have extended this perspective by exploring methods to improve AI models’ ability to express uncertainty (chern2024behonest; yin2023large).
Interactive honesty focuses on the AI model’s ability to maintain objectivity, avoid spreading misinformation, and ensure clarity in its interactions with users. gao2024honestllm define this form of honesty as “the ability to recognize its limitations (e.g., LLMs are unable to process visual information without external tools), remain objective without pandering (e.g., avoiding sycophantic behavior as discussed in (sharma2023towards)), and thereby avoid spreading misinformation or inducing hallucinations.” Interactive honesty emphasizes not only factual accuracy but also the model’s capacity for self-awareness (li2024i; jiang2024assessing) and user-oriented objectivity. This perspective underscores the importance of preventing models from misleading users or offering false reassurance, even in situations where knowledge is incomplete. Subsequent studies have adopted this view to explore honesty through the lens of human cognition and interaction, emphasizing its role in building trust and effective communication (brahman2024art; wen2024know).
@@
Achieving distinct human-centered objectives (labarta2024study) A quantifiable improvement in user performance on tasks that are aligned with the goals facilitated by the provision of explanations.
Peer assessment helpfulness evaluation (liu2024generative) Includes essential features (i.e., comprehensiveness, non-repetitiveness) and constructive features (i.e., praise, problem statement, suggestions, localization, providing examples).
Human-robot collaboration (freedman2020helpfulness) Joint Plan Helpfulness assesses known collaborations, Responsive Plan Helpfulness adapts to dynamic interactions, Normalized Helpfulness standardizes across tasks, and Relative Helpfulness quantifies the reduction in human effort.
Controllable balancing of safety and helpfulness (tuan2024towards) How well the responses fulfill user requests and provide needed information.
Improving LLM honesty and helpfulness simultaneously (gao2024honestllm) The model’s ability to fulfill user requests by providing clear explanations, further guidance, and potential solutions.
The evaluation of AI-generated suggestions in design science research (memmert2023human) Including five dimensions: the ability to inspire new ideas, ease of understanding, relevance to the specific design component, relevance to the broader research domain, and the level of unexpectedness they provide.
RLHF for harmless, honest, and helpful AI (toloka) Understanding the user’s intentions, correctly executing their requested actions, and providing relevant supporting information or alternative solutions if the requested action is not feasible.
@@ 3.3 Case Study III: Helpfulness in Different Contexts
The definition of helpfulness varies substantially based on domain-specific objectives. For instance, in education, hemami2024can frame helpfulness as the ability of AI models to identify optimal learning strategies and personalize content for individual students. For general language assistants, helpfulness is defined as a clear attempt to perform the task or answer the posed question (askell2021general). These variations, illustrated in Table 1, reveal the importance of contextual adaptability in defining and applying the concept of helpfulness.
The absence of standardized definitions and evaluation metrics stems not only from contextual variability but also from researchers’ tendency to optimize for domain-specific objectives. A clear example is the reliance on alignment techniques to tune models to human preferences (ouyang2022training; bai2022training; rafailov2024direct). Most alignment datasets reflect either general value alignment (stiennon2020learning; bai2022constitutional) or specific task objectives (tian2023fine; xie2024v). For example, reasoning-focused tasks prioritize coherent and multi-step explanations (huang2023metatool; zhuo2024astraios), while conversational assistants prioritize concise, user-friendly responses (sun2024adaplanner). Moreover, awad2018moral and santurkar2023whose both reveal that AI models may reflect preference on specific opinions.
Current evaluation methods often employ narrow metrics that align with esearch objectives rather than capturing helpfulness holistically. Moreover, most frameworks utilize specialized metrics (liu2024large) or LLM-as-judge approaches (zheng2023judging; wang2024helpsteer2), ye2024justice, demonstrating that such evaluations can introduce systematic biases through prompt design choices. Future benchmarks should incorporate multidimensional evaluation such as complementary dimensions (e.g., correctness assessment, coherence analysis, creativity metrics), while considering the relative importance and interactions between these aspects. Consequently, multidimensional approaches would better reflect the diverse expectations of helpfulness in real-world applications.
@@ 4 Priority Order
Definition and necessity. In the deployment of AI models, conflicting requirements often arise among helpfulness, honesty, and harmlessness across various scenarios. For instance, in cybersecurity applications, when a user requests information about system vulnerabilities, the model must balance providing helpful technical details (helpfulness) with the need to prevent malicious exploitation (harmlessness) (zhang2024cybench). Similarly, in medical or legal domains, prioritizing honesty may sometimes conflict with harmlessness, such as when conveying critical but distressing information. To address such conflicts, we introduce the concept of priority order, which is defined as: A dynamic hierarchical framework that determines the relative importance and execution sequence of three dimensions of the HHH principle based on contextual requirements.
@@ 4.1 Prioritization Levels and Scales
Prioritization levels refer to the vertical structuring of the HHH principles, meaning that in each task, certain principles take precedence over others based on the risk and ethical constraints of the scenario. As shown in Figure 1, this hierarchical view defines which dimension should be prioritized in different tasks. To characterize the relative importance of each dimension, we introduce the priority value as an attribute representing the extent of emphasis placed on each dimension, which could potentially be quantified through specific metrics in future work. Prior studies have provided empirical evidence supporting the feasibility of prioritization levels across different fields. In education, harmlessness is paramount to protect learners’ development, followed by helpfulness for engagement, with honesty about AI capabilities as the foundation (kooli2023chatbots; selwyn2022future). Creative domains instead prioritize helpfulness to drive innovation, with honesty secondary and harmlessness setting safety bounds while preserving creative freedom (flick2022ethics). This varying prioritization levels across domains indicates the need for an adaptive interpretation of the HHH principle rather than a one-size-fits-all approach.
Prioritization scales refer to horizontal variations within the same ranking level, determining how the HHH principle is applied across different user groups ranging from micro (individual users) to macro (societal user groups). For harmlessness, the micro-scale emphasizes protecting individual privacy and enforcing data minimization (gdpr2016general; staab2024principle), while the macro scale may necessitate selective data collection to mitigate systemic risks and ensure public safety. For instance, an AI system assisting a single user with a research query may prioritize helpfulness and honesty, while an AI generating public information (e.g., automated news summaries) must weigh harmlessness more heavily to prevent misinformation. These scale-dependent variations indicate that effective HHH implementation requires consideration of both individual and societal impacts, even when the nominal priority level remains constant.